# Type Ia supernova growth-rate measurement with LSST simulations: intrinsic scatter systematics

### v1

@@ 4.2 $f\sigma_{8}$ fit results for the four scatter models
Figure 8: Average results over the eight mocks for our four different intrinsic scatter models. The points with error bars represent the mean $f\sigma_{8}$ over the eight mocks, while the wider colored bands represent the mean error obtained over the 8 mocks. The results for true velocities are in green, the results of the simple fit are in blue and the results of the BBC fit are in red.
@@
In Fig. 8 and Table 4 we present the results of the fitting procedure for the four intrinsic scatter models. Fitting for $f\sigma_{8}$ with the true velocities, which correspond to a test case, gives unbiased results for all models with a mean error of about $\sim 5\%$ that represents the statistical limit of our simulated data. The fact that this fit is unbiased is a hint that the particular distribution of SN Ia host galaxy masses does not lead to a particular velocity statistic that could bias the $f\sigma_{8}$ measurement. For the $S_{\mathrm{COH}}$ simulation, as expected, the simple and the BBC fitting procedure perform largely the same with an unbiased value and an error of $\sim 14\%$. For the $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$ models, the two fitting procedures perform also very similarly with error on $f\sigma_{8}$ of about $13\%$. However, for the $S_{\mathrm{C11}}$ simulation, a bias of $\sim 4\%$ with a significance of $\sim 2\sigma$ seems to be present in both the simple and the BBC fit. This bias is negligible when compared to the $13\%$ averaged uncertainty. For our most realistic simulation, $S_{\mathrm{P23}}$, the results for the two fitting methods are biased with a strong significance. The simple method is slightly less biased, with a bias of $\sim-20\%$ against a bias of $\sim-26\%$ for the BBC fit. We understand this bias as a consequence of the non-Gaussianity of the estimated velocities noted in the previous section. For the mean error on $f\sigma_{8}$ we obtain $\sim 11\%$ from the BBC fit and $\sim 15\%$ from the simple fit.
@@
Table 4 shows also that, for the $S_{\mathrm{COH}}$, $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$, the mean errors on $f\sigma_{8}$ obtained from the fit seem to be slightly higher than the scattering of the measures computed over the eight mocks. This could point to an overestimation of the errors that should be investigated in future works. This effect does not appear for the $S_{\mathrm{G10}}$ simulation, which shows the opposite behavior, with a mean error smaller than the scatter. We note that while the scattering of the results over our eight mocks is similar for the simple and the BBC fit in the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations with values of $\sim 7.5\%$ for $S_{\mathrm{COH}}$ and $\sim 16\%$ for $S_{\mathrm{G10}}$. In the case of the most chromatic models $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$ the scatter is less important for the BBC fit.For $S_{\mathrm{C11}}$ we find a scatter of the results of about $\sim 7\%$ for the simple fit and of about $\sim 6\%$ for the BBC fit, for $S_{\mathrm{P23}}$ we find a scatter of $\sim 12\%$ for the simple fit and $\sim 9\%$ for the BBC fit. We note that these results are using the low statistics of only eight realizations and may also be subject to the small correlations existing between the sub-mocks.
@@ 4.3 Propagation of the uncertainty on the BS21 intrinsic scatter model parameters to $f\sigma_{8}$
Here, we test the impact of adding the covariance matrix that represents the uncertainty on the BS21 model parameters. We compute this covariance matrix in the same way as stated in Sect. 3.2 of V24. During the construction of the BS21 intrinsic scatter model covariance matrix, an extra cut is added on the SN Ia sample by BBC. This cut requires that the corrected distance modulus passes the cuts presented in Sect. 3.3 for each of the bias corrections obtained from the different bias correction simulations. This cut slightly reduces our sample from $\langle N_{\mathrm{SN}}\rangle\sim 6916$ to $\langle N_{\mathrm{SN}}\rangle\sim 6630$.
The fit results are presented in Table 5 as an average over the eight mocks. The fit of the true velocities gives, as before, unbiased results with a relative error of 4.9%. The sub-sampling due to the additional cut slightly worsens the simple fit bias to a value of $\sim-29\%$ with the a slightly higher relative error of $\sim 16\%$. The bias of the BBC fit without the covariance matrix is slightly increased by the sub-sampling with a value of $\sim-35\%$ and a slightly higher relative error of $\sim 12\%$. The BBC fit that includes the intrinsic scatter model covariance matrix does not show any significant variations from the one without the inclusion of the covariance matrix, with an average shift of $\sim-0.2\%$ and no changes on the $f\sigma_{8}$ error. From this we conclude that the error on $f\sigma_{8}$ is not dominated by the realizations of the P23 scatter model parameters, as when measuring $w$, but rather the amplitude of the Hubble residual scatter. This difference with respect to other cosmological analyses such as the constraints on the dark energy equation of state can be explained by the fact that in the $f\sigma_{8}$ analysis the statistical error is scaled down by the density of SNe Ia and not by their total number. Hence, the intrinsic scatter model systematic does not contribute significantly to the error budget for $f\sigma_{8}$.
Table 5: Results for the $S_{\mathrm{P23}}$ simulation with and without including the intrinsic scatter covariance matrix from BS21 parameter uncertainty. The first column shows the average of $f\sigma_{8}$ fitted values and the second column show the average uncertainty on $f\sigma_{8}$ obtain over our eight mocks.
@@ 4.4 The $\sigma_{u}$ systematic
Using the average of the result of the $f\sigma_{8}$ fit with the true velocity over the eight mocks, we find that the range of $\sigma_{u}$ that allows us to recover the fiducial value of $f\sigma_{8}$ in a $5\sigma$ confidence range is $\sigma_{u}\in\left[19,24\right]$ Mpc $h^{-1}$. We note that the best-fit $\sigma_{u}$ values found with the fit on true velocities is $\sigma_{u}\simeq 21.5$ Mpc $h^{-1}$, slightly higher than the value we calibrated in Sect. 3.4.2 and used as a baseline, $\sigma_{u}\simeq 21$ Mpc $h^{-1}$. The range we found is consistent with the one found in Table 1 of Y. Lai et al. (2023) that is $\sigma_{u}\in\left[19,23\right]$ Mpc $h^{-1}$. By applying Eq. 47 on this range we find $\sigma_{f\sigma_{8}}^{\sigma_{u}}\sim 6\%$, contributing to roughly $\sim$23of the total error budget of $f\sigma_{8}$. This result is again similar to the variation shown in Table 1 of Y. Lai et al. (2023) in which $f\sigma_{8}$ varies by $\sim 7\%$ over a $\sigma_{u}$ range between 19 Mpc $h^{-1}$ and 23 Mpc $h^{-1}$.
@@ 5 Conclusion
In this paper, we simulated eight realizations of the 10 years of SN Ia observations from the WFD program of the LSST survey for four different intrinsic scatter models. Our simulations include realistic distribution of SN Ia host masses and correlations between SN Ia SALT parameters and hosts. We build the SN Ia Hubble diagram and estimate velocities, comparing two different approaches. The simple method, used in C23, is a naive construction of the Hubble diagram from the Tripp relation. The second, the BBC approach, uses large simulations to correct for biases in the Hubble diagram. We show that the two methods have comparable estimates of PVs for the two less chromatic intrinsic scatter models. However, we find that for the two more chromatic models the simple approach slightly biases the velocities with respect to the color parameter, $c$. The BBC approach corrects this bias in PV estimates.
@@
Furthermore, we find that including the covariance due to the uncertainty on the BS21 model parameters does not significantly increase the error on $f\sigma_{8}$, with a negligible contribution to the total error budget. Overall, the error budget of $f\sigma_{8}$ is dominated by the statistical error that represents $>75\%$ of the total error on $f\sigma_{8}$. We evaluate that the current systematic error is dominated by the uncertainty on the value of the nuisance parameter of the damping function $\sigma_{u}$. We quantify that the uncertainty on $\sigma_{u}$ results in a systematic error on $f\sigma_{8}$ of about $\sim 6$%. This last result further motivates the ongoing effort to find a more physically motivated parametrization of the redshift space distortion effect on the velocity power spectrum (L. Dam et al., 2021).
This paper has undergone internal review in the LSST Dark Energy Science Collaboration. The internal reviewers were Dillon Brout and Tamara M. Davis. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy and the LSST Discovery Alliance in the United States. DESC uses resources of the IN2P3 Computing Center (CC-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. J.E.B. acknowledges funding from Excellence Initiative of Aix-Marseille University - A*MIDEX, a French “Investissements d’Avenir” program (AMX-20-CE-02 - DARKUNI). L.G. acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, CEX2020-001058-M, ILINK23001, COOPB2304, and 2021-SGR-01270. {contribution} BC performed the simulations, the analysis, and wrote the majority of the paper. R. C. contributed to the simulation setup, paper writing and review and participated in analysis discussions. E. R. P. contributed to the writing and review of the paper and participated in analysis discussions. D. S. contributed to the writing and review of the paper and participated in analysis discussions. C. R. contributed to the code development and review. D. R. reviewed the code and the manuscript.
@@ Appendix A Fit results for the SNe Ia standardization parameters
For the color correction linear coefficient $\beta$, the fitted values for the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations are slightly underestimated when using the simple method, while the values obtained with the BBC method are compatible with the input value of $\beta=3.1$ within the averaged uncertainty. In the case of the $S_{\mathrm{C11}}$ simulation the simple method gives an underestimated value when compared to the input value of $\beta=3.8$, while the BBC method value is compatible with the input value within the averaged uncertainty. For the $S_{\mathrm{P23}}$ simulation, there is no input value of $\beta$; however the value recovered by both the simple and the BBC method seem compatible and are similar to the value of $\beta\sim 2.83$ found on DES-SN+low-z simulations in Table 8 of V24.
Finally, for the mass step parameter $\gamma$, we find that for the $S_{\mathrm{COH}}$, $S_{\mathrm{G20}}$ and $S_{\mathrm{C11}}$ simulations, the simple fit recovers the value of the input $\gamma=0.050$ within the average uncertainty error. The BBC fit recovers the value of $\gamma$ for the $S_{\mathrm{C11}}$ simulation but presents a bias of $\sim 1.5\sigma$ with respect to the averaged uncertainty for the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations. In the case of the $S_{\mathrm{P23}}$ there is no input value for $\gamma$ since the mass-step emerges from the dust extinction-host mass correlation. The value of $\gamma$ fitted using the BBC mehtod is close to a value of $\gamma\sim 0$, a result similar to what was found on DES simulations in Table 4 of V24.

### v2

@@ 4.2 $f\sigma_{8}$ fit results for the four scatter models
Figure 8: Average results over the eight mocks for our four different intrinsic scatter models. The points with error bars represent the mean $f\sigma_{8}$ over the eight mocks, while the wider colored bands represent the mean error obtained over the 8 mocks. The results for true velocities are in green, the results of the simple fit are in blue, and the results of the BBC fit are in red.
@@
In Fig. 8 and Table 4, we present the results of the fitting procedure for the four intrinsic scatter models. Fitting for $f\sigma_{8}$ with the true velocities, which correspond to a test case, gives unbiased results for all models with a mean error of about $\sim 5\%$ that represents the statistical limit of our simulated data. The fact that this fit is unbiased is a hint that the particular distribution of SN Ia host galaxy masses does not lead to a particular velocity statistic that could bias the $f\sigma_{8}$ measurement. For the $S_{\mathrm{COH}}$ simulation, as expected, the simple and the BBC fitting procedure perform largely the same with an unbiased value and an error of $\sim 14\%$. For the $S_{\mathrm{G10}}$ and $S_{\mathrm{C11}}$ models, the two fitting procedures perform also very similarly with error on $f\sigma_{8}$ of about $13\%$. However, for the $S_{\mathrm{C11}}$ simulation, a bias of $\sim 4\%$ with a significance of $\sim 2\sigma$ seems to be present in both the simple and the BBC fit. This bias is negligible when compared to the $13\%$ averaged uncertainty. For our most realistic simulation, $S_{\mathrm{P23}}$, the results for the two fitting methods are biased with a strong significance. The simple method is slightly less biased, with a bias of $\sim-20\%$ against a bias of $\sim-26\%$ for the BBC fit. We understand this bias as a consequence of the non-Gaussianity of the estimated velocities noted in the previous section. For the mean error on $f\sigma_{8}$ we obtain $\sim 11\%$ from the BBC fit and $\sim 15\%$ from the simple fit.
@@
Table 4 shows also that, for the $S_{\mathrm{COH}}$, $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$, the mean errors on $f\sigma_{8}$ obtained from the fit seem to be slightly higher than the scattering of the measures computed over the eight mocks. This could point to an overestimation of the errors that should be investigated in future works. This effect does not appear for the $S_{\mathrm{G10}}$ simulation, which shows the opposite behavior, with a mean error smaller than the scatter. We note that while the scattering of the results over our eight mocks is similar for the simple and the BBC fit in the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations, with values of $\sim 7.5\%$ for $S_{\mathrm{COH}}$ and $\sim 16\%$ for $S_{\mathrm{G10}}$. In the case of the most chromatic models $S_{\mathrm{C11}}$ and $S_{\mathrm{P23}}$ the scatter is less important for the BBC fit.For $S_{\mathrm{C11}}$ we find a scatter of the results of about $\sim 7\%$ for the simple fit and of about $\sim 6\%$ for the BBC fit, for $S_{\mathrm{P23}}$ we find a scatter of $\sim 12\%$ for the simple fit and $\sim 9\%$ for the BBC fit. We note that these results are using the low statistics of only eight realizations and may also be subject to the small correlations existing between the sub-mocks.
@@ 4.3 Propagation of the uncertainty on the BS21 intrinsic scatter model parameters to $f\sigma_{8}$
Here, we test the impact of adding the covariance matrix that represents the uncertainty on the brout_its_2021 model parameters. We compute this covariance matrix in the same way as stated in Sect. 3.2 of vincenzi_dark_2024. During the construction of the brout_its_2021 intrinsic scatter model covariance matrix, an extra cut is added on the SN Ia sample by BBC. This cut requires that the corrected distance modulus passes the cuts presented in Sect. 3.3 for each of the bias corrections obtained from the different bias correction simulations. This cut slightly reduces our sample from $\langle N_{\mathrm{SN}}\rangle\sim 6916$ to $\langle N_{\mathrm{SN}}\rangle\sim 6630$.
The fit results are presented in Table 5 as an average over the eight mocks. The fit of the true velocities gives, as before, unbiased results with a relative error of 4.9%. The sub-sampling due to the additional cut slightly worsens the simple fit bias to a value of $\sim-29\%$ with the a slightly higher relative error of $\sim 16\%$. The bias of the BBC fit without the covariance matrix is slightly increased by the sub-sampling with a value of $\sim-35\%$ and a slightly higher relative error of $\sim 12\%$. The BBC fit that includes the intrinsic scatter model covariance matrix does not show any significant variations from the one without the inclusion of the covariance matrix, with an average shift of $\sim-0.2\%$ and no changes on the $f\sigma_{8}$ error. From this, we conclude that the error on $f\sigma_{8}$ is not dominated by the realizations of the P23 scatter model parameters, as when measuring $w$, but rather the amplitude of the Hubble residual scatter. This difference with respect to other cosmological analyses such as the constraints on the dark energy equation of state can be explained by the fact that in the $f\sigma_{8}$ analysis the statistical error is scaled down by the density of SNe Ia and not by their total number. Hence, the intrinsic scatter model systematic does not contribute significantly to the error budget for $f\sigma_{8}$.
Table 5: Results for the $S_{\mathrm{P23}}$ simulation with and without including the intrinsic scatter covariance matrix from brout_its_2021 parameter uncertainty. The first column shows the average of $f\sigma_{8}$ fitted values and the second column show the average uncertainty on $f\sigma_{8}$ obtain over our eight mocks.
@@ 4.4 The $\sigma_{u}$ systematic
Using the average of the result of the $f\sigma_{8}$ fit with the true velocity over the eight mocks, we find that the range of $\sigma_{u}$ that allows us to recover the fiducial value of $f\sigma_{8}$ in a $5\sigma$ confidence range is $\sigma_{u}\in\left[19,24\right]$ Mpc $h^{-1}$. We note that the best-fit $\sigma_{u}$ values found with the fit on true velocities is $\sigma_{u}\simeq 21.5$ Mpc $h^{-1}$, slightly higher than the value we calibrated in Sect. 3.4.2 and used as a baseline, $\sigma_{u}\simeq 21$ Mpc $h^{-1}$. The range we found is consistent with the one found in Table 1 of lai_using_2023 that is $\sigma_{u}\in\left[19,23\right]$ Mpc $h^{-1}$. By applying Eq. 47 on this range we find $\sigma_{f\sigma_{8}}^{\sigma_{u}}\sim 6\%$, contributing to roughly $\sim$23% of the total error budget of $f\sigma_{8}$. This result is again similar to the variation shown in Table 1 of lai_using_2023 in which $f\sigma_{8}$ varies by $\sim 7\%$ over a $\sigma_{u}$ range between 19 Mpc $h^{-1}$ and 23 Mpc $h^{-1}$.
@@ 5 Conclusion
In this paper, we simulated eight realizations of the 10 years of SN Ia observations from the WFD program of the LSST survey for four different intrinsic scatter models. Our simulations include realistic distribution of SN Ia host masses and correlations between SN Ia SALT parameters and hosts. We build the SN Ia Hubble diagram and estimate velocities, comparing two different approaches. The simple method, used in carreres_growth-rate_2023, is a naive construction of the Hubble diagram from the Tripp relation. The second, the BBC approach, uses large simulations to correct for biases in the Hubble diagram. We show that the two methods have comparable estimates of PVs for the two less chromatic intrinsic scatter models. However, we find that for the two more chromatic models the simple approach slightly biases the velocities with respect to the color parameter, $c$. The BBC approach corrects this bias in PV estimates.
@@
Furthermore, we find that including the covariance due to the uncertainty on the brout_its_2021 model parameters does not significantly increase the error on $f\sigma_{8}$, with a negligible contribution to the total error budget. Overall, the error budget of $f\sigma_{8}$ is dominated by the statistical error that represents $75\%$ of the total error on $f\sigma_{8}$. We evaluate that the current systematic error is dominated by the uncertainty on the value of the nuisance parameter of the damping function $\sigma_{u}$. We quantify that the uncertainty on $\sigma_{u}$ results in a systematic error on $f\sigma_{8}$ of about $\sim 6$%. This last result further motivates the ongoing effort to find a more physically motivated parametrization of the redshift space distortion effect on the velocity power spectrum (dam_exploring_2021).
This paper has undergone internal review in the LSST Dark Energy Science Collaboration. The internal reviewers were Dillon Brout and Tamara M. Davis. The DESC acknowledges ongoing support from the Institut National de Physique Nucléaire et de Physique des Particules in France; the Science & Technology Facilities Council in the United Kingdom; and the Department of Energy and the LSST Discovery Alliance in the United States. DESC uses resources of the IN2P3 Computing Center ($C$-IN2P3–Lyon/Villeurbanne - France) funded by the Centre National de la Recherche Scientifique; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231; STFC DiRAC HPC Facilities, funded by UK BEIS National E-infrastructure capital grants; and the UK particle physics grid, supported by the GridPP Collaboration. This work was performed in part under DOE Contract DE-AC02-76SF00515. J.E.B. acknowledges funding from Excellence Initiative of Aix-Marseille University - A*MIDEX, a French “Investissements d’Avenir” program (AMX-20-CE-02 - DARKUNI). L.G. acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, CEX2020-001058-M, ILINK23001, COOPB2304, and 2021-SGR-01270. {contribution}B. C. performed the simulations, the analysis, and wrote the majority of the paper. R. C. contributed to the simulation setup, paper writing and review, and participated in analysis discussions. E. R. P. contributed to the writing and review of the paper and participated in analysis discussions. D. S. contributed to the writing and review of the paper and participated in analysis discussions. C. R. contributed to the code development and review. D. R. reviewed the code and the manuscript.
@@ Appendix A Fit results for the SNe Ia standardization parameters
For the color correction linear coefficient $\beta$, the fitted values for the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations are slightly underestimated when using the simple method, while the values obtained with the BBC method are compatible with the input value of $\beta=3.1$ within the averaged uncertainty. In the case of the $S_{\mathrm{C11}}$ simulation the simple method gives an underestimated value when compared to the input value of $\beta=3.8$, while the BBC method value is compatible with the input value within the averaged uncertainty. For the $S_{\mathrm{P23}}$ simulation, there is no input value of $\beta$; however the value recovered by both the simple and the BBC method seem compatible and are similar to the value of $\beta\sim 2.83$ found on DES-SN+low-z simulations in Table 8 of vincenzi_dark_2024.
Finally, for the mass step parameter $\gamma$, we find that for the $S_{\mathrm{COH}}$, $S_{\mathrm{G20}}$ and $S_{\mathrm{C11}}$ simulations, the simple fit recovers the value of the input $\gamma=0.050$ within the average uncertainty error. The BBC fit recovers the value of $\gamma$ for the $S_{\mathrm{C11}}$ simulation but presents a bias of $\sim 1.5\sigma$ with respect to the averaged uncertainty for the $S_{\mathrm{COH}}$ and $S_{\mathrm{G10}}$ simulations. In the case of the $S_{\mathrm{P23}}$ there is no input value for $\gamma$ since the mass-step emerges from the dust extinction-host mass correlation. The value of $\gamma$ fitted using the BBC method is close to a value of $\gamma\sim 0$, a result similar to what was found in DES simulations in Table 4 of vincenzi_dark_2024.
@@
Appendix B Non-Gaussianties and variations of the P23 parameters
To further understand the origin of non-Gaussianities arising from the P23 dust model, we investigate the consequences of varying some parameters of the model. In particular, we simulate our sample with a reduced variance $\tau_{E_{\mathrm{dust}}}$ of the dust-extinction exponential distribution. We simulated the SNe Ia sample for different values of $\tau_{E_{\mathrm{dust}}}$ reduced by 25, 50, 75, 95, and 100% with respect to the value used in our nominal $S_{\mathrm{P23}}$ simulation. In Fig. 10 we show the color – velocity-bias density contours for these different simulations, compared to that of $S_{\mathrm{P23}}$. We can see that the color bias and the asymmetry of the velocity distribution are growing along with the increase of the $\tau_{\mathrm{E_{\mathrm{dust}}}}$ parameter.
Figure 10: Bias in estimated velocity as a function of fitted SALT color parameter $c$. The contours represent the 25%, 50% and 75% density levels of the SN Ia sample and points are the velocity bias in different $c$ bins. The blue contours correspond to $S_{\mathrm{P23}}$, while yellow contours are obtained using variation of the $\tau_{\mathrm{E_{dust}}}$ parameter.
We fitted for $f\sigma_{8}$ with these simulations using the simple fitting method. We did not run the BBC fitting method since it would require simulating a bias correction for each variation of $\tau_{E_{\mathrm{dust}}}$. For each mock, simulations of variation of $\tau_{E_{\mathrm{dust}}}$ use the same random seed. Thus, the host galaxies and velocities used in the fit of $f\sigma_{8}$ are the same and the differences lie in the fitted SALT parameters. The results for $f\sigma_{8}/(f\sigma_{8})_{\mathrm{fid}}$ are shown in Fig. 11 as a function of the ratio between the simulated $\tau_{E_{\mathrm{dust}}}$ and our baseline value used in $S_{\mathrm{P23}}$. We see that the increase of $\tau_{E_{\mathrm{dust}}}$ is correlated with a decrease of the bias in the fitted value of $f\sigma_{8}$. In the “No dust” case ($\tau_{E_{\mathrm{dust}}}=0$) we remark that we still have a bias of $\sim-9\%$ with respect to the fiducial $f\sigma_{8}$ value. Since the Hubble scatter is decreasing with the suppression of dust, this remaining bias could be due to other systematic effects that dominate this unrealistic low-scatter regime, and we let there identifications to later studies. We also tested reducing the scatter $\sigma_{\beta}$ of the distribution of $\beta$ but we found its impact on the $f\sigma_{8}$ bias to be negligible.
Figure 11: Average result over the eight mocks as a function of variations of the value of the $\tau_{E_{\mathrm{dust}}}$ with respect to the baseline value used in $S_{\mathrm{P23}}$.